Would someone mind helping me on these please?
The perimeter of a rectangle is 74 m. The length is 7 m more than twice the width. Find the dimensions.
What is the length?
Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 53 points. They won 10 more games than they tied. How many wins and how many ties did the team have?

The perimeter of a rectangle is 158 inches. The length exceeds the width by 55 inches. Find the length and the width.

The Everton College store paid $1270 for an order of 28 calculators. The store paid $9 for each scientific calculator. The others, all graphing calculators, cost the store $55 each. How many of each type of calculator was ordered?

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

The perimeter of a rectangle is 74 m. The length is 7 m more than twice the width. Find the dimensions.
What is the length?
2L+2W=74
L=2w+7
The perimeter of a rectangle is 158 inches. The length exceeds the width by 55 inches. Find the length and the width.
2L+2W=158
L=W+55
Try to submit one problem per submission.

Answer by stanbon(75887)

The perimeter of a rectangle is 74 m. The length is 7 m more than twice the width. Find the dimensions.
What is the length?
---
Equation:
Perimeter = 2(Length + Width)
74 = 2((2W+7) + W)
37 = 3W = 7
3W = 30
W = 10 m (width)
Length = L = 2W+7 = 2*30+7 = 67 m (length)
-------------------------------------------------

Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 53 points. They won 10 more games than they tied. How many wins and how many ties did the team have?
Equations:
Quantity Equation: W = T + 10
Points Equation: 2W + T = 53
-------------------
Substitute for W and solve for "T":
2(T+10) + T = 53
3T = 33
T = 11 (# of ties)
---
Since W = T + 10, W = 11+10 = 21 (# of wins)
==================================================
The perimeter of a rectangle is 158 inches. The length exceeds the width by 55 inches. Find the length and the width.
---
Follow the method in # 1 above.
------
The Everton College store paid $1270 for an order of 28 calculators. The store paid $9 for each scientific calculator. The others, all graphing calculators, cost the store $55 each. How many of each type of calculator was ordered?
---
Equations:
Value Eq: 9S + 55G2 = 1270
Quantity Eq: S + G = 28
---------------------------
S = 28 - G
---
Substitute for S and solve for "C":
9(28-G) + 55G = 1270
252 -9G + 55G = 1270
46G = 1018
G = 22.13..
---
Comment: The value should be a whole number
as G is a count of graphing calculators.
Check you posted values.
-----
Cheers,
Stan H.